OpenStudy (anonymous):

3t – 18t + 4 = -26 please show all work

OpenStudy (anonymous):

3t – 18t + 4 = -26 -15t+4=-26 -4 -4 -15t=-30 -15t/-15=-30/-15 t=2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem